# Tensor Gaussian Processes: Efficient Solvers for Nonlinear PDEs

### v1

@@ Abstract
Machine learning solvers for partial differential equations (PDEs) have attracted growing interest. However, most existing approaches, such as neural network solvers, rely on stochastic training, which is inefficient and typically requires a great many training epochs. Gaussian process (GP)/kernel-based solvers, while mathematical principled, suffer from scalability issues when handling large numbers of collocation points often needed for challenging or higher-dimensional PDEs. To overcome these limitations, we propose TGPS, a tensor-GP-based solver that models factor functions along each input dimension using one-dimensional GPs and combines them via tensor decomposition to approximate the full solution. This design reduces the task to learning a collection of one-dimensional GPs, substantially lowering computational complexity, and enabling scalability to massive collocation sets. For efficient nonlinear PDE solving, we use a partial freezing strategy and Newton’s method to linerize the nonlinear terms. We then develop an alternating least squares (ALS) approach that admits closed-form updates, thereby substantially enhancing the training efficiency. We establish theoretical guarantees on the expressivity of our model, together with convergence proof and error analysis under standard regularity assumptions. Experiments on several benchmark PDEs demonstrate that our method achieves superior accuracy and efficiency compared to existing approaches.
@@ 1 Introduction
Machine learning (ML) solvers for partial differential equations (PDEs) have been receiving increasing attention due to their ease of implementation and competitive accuracy. These approaches approximate the solution with a machine learning model, such as deep neural networks (Raissi et al., 2019a, ), trained by minimizing a composite objective function that combines boundary and residual losses evaluated at a set of collocation points, thereby enforcing the boundary conditions and equation. Unlike traditional numerical solvers, ML approaches avoid complex, problem-specific discretization schemes and numerical routines, making them simpler and more convenient to implement and verify.
Despite these advantages, most existing ML solvers — including physics-informed neural networks (Raissi et al., 2019b, ) and recent Gaussian process (GP) and kernel-based methods (Fang et al.,, 2023; Xu et al.,, 2024) — rely on stochastic optimization to effectively learn the model parameters, which often requires tens of thousands to even millions of iterations, making solving procedure quite inefficient. GP and kernel-based solvers, though mathematically principled, also face scalability challenges as the number of collocation points grows. For example, Chen et al., 2021a ; Long et al., (2022) place a GP prior over the solution and its derivatives, yielding block-structured covariance matrices whose time and memory costs exceed the standard $\mathcal{O}(M^{3})$ and $\mathcal{O}(M^{2})$ scaling, where $M$ is the number of collocation points. To mitigate this, Fang et al., (2023); Xu et al., (2025) proposed using product kernels on Cartesian grids, exploiting Kronecker algebra for efficient matrix operations. However, this approach requires estimating the solution values at all grid points, leading to exponential growth in parameters with dimension, and its reliance on structured grids limits applicability to irregular domains.
@@
Model: We introduce a set of one-dimensional factor functions, each modeled as a GP, along every input dimension. These factors are combined via multilinear tensor decompositions — such as CANDECOMP/PARAFAC (CP) (Harshman et al.,, 1970) or tensor-ring decomposition (Zhao et al.,, 2016) — to approximate the full solution. For each dimension, we place inducing points and represent the factor function by its GP conditional mean, with the inducing values serving as trainable parameters. Collocation points are freely sampled to form the training objective. This design allows our model to scale linearly with both the PDE dimension and the number of collocation points — not only in covariance matrix computation and storage but also in the number of trainable parameters.
@@ 2 Background
where ${\mathcal{F}}_{r}=\frac{1}{M_{\Omega}}\sum_{m=1}^{M_{\Omega}}\left({\mathcal{P}}(\text{NN}_{\Theta})({\bf x}_{m})-a({\bf x}_{m})\right)^{2}$ is the PDE residual loss, and ${\mathcal{F}}_{b}=\frac{1}{M-M_{\Omega}}\sum_{j=1}^{M-M_{\Omega}}\left({\mathcal{B}}\left(\text{NN}_{\Theta}\right)({\bf x}_{M_{\Omega}+j})-b({\bf x}_{M_{\Omega}+j})\right)^{2}$ is the boundary loss. Here NN denotes the neural network, $\Theta$ its parameters, and $\lambda_{b}>0$ a weighting coefficient. Training typically relies on stochastic optimization (e.g., ADAM (Kingma,, 2014)), as in standard neural network applications like image classification. Achieving good accuracy generally requires tens of thousands of iterations, and a second-order optimizer (e.g., L-BFGS) is often used for refinement and stabilization (Shin,, 2020; Li et al.,, 2023; Penwarden et al.,, 2023). As a result, while this framework is straightforward and convenient to implement, the training (solving) process is often lengthy and inefficient.
@@
Gaussian Process and Kernel-Based Solvers. An alternative class of solvers is based on GP/kernel methods, which rest on strong mathematical foundations. In (Chen et al., 2021a, ; Long et al.,, 2022) , a GP prior is placed over the solution $u$ and all its derivatives (or more generally, linear operators) $D_{j}(u)$ appearing in the PDE. The goal is to estimate the values of $u$ and all $D_{j}(u)$ evaluated at the collocation points, which leads to a multi-variate Gaussian prior distribution with a block covariance matrix, ${\bf C}=\{{\bf C}_{ij}\}$, where each block ${\bf C}_{ij}$ is associated with a pair of linear operators (e.g., derivatives and $u$ itself). Since each collocation point can contribute to multiple values (e.g., different $D_{j}(u)$), the covariance matrix $\bf C$} is typically larger than $M\times M$, where $M$ is the number of collocation points. As $M$ increases, the time and space complexity exceed $\mathcal{O}(M^{3})$ and $\mathcal{O}(M^{2})$, respectively, making computation prohibitively expensive or even infeasible for large $M$.
@@
To mitigate this issue, recent work by Xu et al., (2025) approximates the solution using standard GP/kernel interpolation: $u({\bf x};{\bf u}_{{\mathcal{M}}})=\kappa({\bf x},{\mathcal{M}}){\bf K}_{MM}^{-1}{\bf u}_{{\mathcal{M}}}$, where ${\bf K}_{MM}=\kappa({\mathcal{M}},{\mathcal{M}})$ is the $M\times M$ covariance matrix at the collocation points, and $\bf u$$_{{\mathcal{M}}}$ denotes the solution values at these points. Differential operators $D_{j}$ are applied directly to this interpolation to approximate $D_{j}(u)$, enabling the use of a training loss similar to (2). To reduce computation cost, collocation points are placed on a Cartesian grid: ${\mathcal{M}}={\bf s}^{1}\times\ldots\times{\bf s}^{d}$ where each ${\bf s}^{j}$ is a set of input locations along dimension $j$. By adopting a product kernel of the form $\kappa(\mathbf{x},\mathbf{x}^{\prime})=\prod_{j}\kappa_{j}(x_{j},x_{j}^{\prime})$, the covariance matrix admits a Kronecker product structure: $\kappa({\mathcal{M}},{\mathcal{M}})={\bf K}_{1}\otimes\ldots\otimes{\bf K}_{d}$, where each ${\bf K}_{j}=\kappa_{j}({\bf s}^{j},{\bf s}^{j})$ is the kernel matrix along input dimension $j$. Kronecker algebra (Kolda,, 2006) allows one to avoid computing the full ${\bf K}_{MM}$; instead, only the smaller matrices ${\bf K}_{j}$ are computed and inverted, greatly reducing the cost. However, this method still requires estimating $\mathcal{U}$$_{{\mathcal{M}}}$ — the solution values over the entire grid $\mathcal{M}$}. As the input dimension increases, the size of $\mathcal{U}$$_{\mathcal{M}}$ grows exponentially, making storage and estimation infeasible. Furthermore, effective training still relies on stochastic optimization, often requiring up to one million iterations (Xu et al.,, 2025), rendering the solving procedure inefficient in practice.
@@ 3.1 Model
To leverage the principled mathematical framework of GPs/kernel methods while overcoming their scalability and efficiency bottlenecks, we propose TGPS, a tensor-GP based PDE solver. Specifically, we introduce a set of one-dimensional factor functions for each input dimension $i$,
@@
where $1\leq i\leq d$, and ${\mathcal{G}}^{i}$ is a Reproducing Kernel Hilbert Space (RKHS) induced by a Mercer kernel function $\kappa_{i}(\cdot,\cdot)$. In other words, we model each factor function as a GP, $f^{i}_{r}\sim\mathcal{GP}(0,\kappa_{i}(\cdot,\cdot))$. Without loss of generality, we assume the PDE domain $\Omega\cup\partial\Omega\subseteq\bigtimes_{i=1}^{d}\Omega_{0}$. We then combine these factor functions via mutilinear tensor decomposition to construct the solution approximation. We consider two representative tensor decomposition models.
CANDECOMP/PARAFAC(CP) Decomposition (Harshman et al.,, 1970). We set $R_{1}=\ldots=R_{d}=R$, and model the solution function as
@@
Tensor-Ring (TR) Decomposition (Zhao et al.,, 2016). In each dimension $i$, we view the factor functions together as a single function with a matrix output, ${\bf F}^{i}=\{f^{i}_{r}(\cdot)\}:\mathbb{R}\rightarrow\mathbb{R}^{R_{i-1}\times R_{i}}$ where $R_{0}=R_{d}$. The solution is modeled as
@@ 3.2 Algorithm
Newton’s method. Consider solving the PDE as a root finding problem: $\mathcal{R}(u)=0$, where $\mathcal{R}$} is the PDE operator. We apply Newton’s method by linearizing ${\mathcal{R}}(u)$ around the previous iteration $u^{\text{prev}}$ via a first-order Taylor expansion: ${\mathcal{R}}(u)\approx{\mathcal{R}}(u^{\text{prev}})+J(u^{\text{prev}})(u-u^{\text{prev}})=0$, where $J(u^{\text{prev}})$ is the Fréchet derivative, $J(u^{\text{prev}})=\frac{{\rm d}\mathcal{R}(u^{\text{prev}}+\epsilon v)}{{\rm d}\epsilon}\bigg|_{\epsilon=0}$, and $v$ is an arbitrarily small perturbation. By construction, this first-order approximation is always linear in $u$. In practice, the procedure amounts to replacing the nonlinear terms in the PDE and boundary conditions with their first-order Taylor expansions. For example, in (9), the cubic term $u^{3}$ is replaced by $(u^{\text{prev}})^{3}+3(u^{\text{prev}})^{2}(u-u^{\text{prev}})$.
@@ 4 Theoretical Analysis
Next, we establish the convergence result under the CP format. Our analysis follows the roadmap of Batlle et al., (2023); Xu et al., (2025), adopting standard assumptions on PDE stability and on the regularity of the domain and boundary conditions (Xu et al.,, 2025, Assumption 4.1).
@@
Let $u^{*}\in{\mathcal{U}}$ denote the unique strong solution of (1), and suppose Assumption 4.3 holds. Let ${\mathcal{M}}={\mathcal{M}}_{\Omega}\cup{\mathcal{M}}_{\partial\Omega}$ be a set of collocation points, with ${\mathcal{M}}_{\Omega}\subset\Omega$ and ${\mathcal{M}}_{\partial\Omega}\subset\partial\Omega$. Assume the Voronoi diagram induced by $\mathcal{M}$} has a uniformly bounded aspect ratio across all cells. Define the fill-distances $h_{\Omega}:=\underset{{\bf x}\in\Omega}{\sup}\;\underset{{\bf x}^{\prime}\in{\mathcal{M}}_{\Omega}}{\inf}|{\bf x}-{\bf x}^{\prime}|$ and $h_{\partial\Omega}:=\underset{{\bf x}\in\partial\Omega}{\sup}\;\;\underset{{\bf x}^{\prime}\in{\mathcal{M}}_{\partial\Omega}}{\inf}\rho_{\partial\Omega}({\bf x},{\bf x}^{\prime})$, where $|\cdot|$ is the Euclidean distance, and $\rho_{\partial\Omega}$ is a geodesic distance defined on $\partial\Omega$. Set $h=\text{max}(h_{\Omega},h_{\partial\Omega})$. Suppose each RKHS ${\mathcal{G}}^{i}$, to which the factor functions in dimension $i$ belong, is associated with a universal kernel, and that ${\mathcal{U}}={\mathcal{G}}^{1}\otimes\cdots\otimes{\mathcal{G}}^{d}$. Then, for any arbitrarily small $\varepsilon>0$, with a sufficiently large $R$ and an appropriate $\delta$, the optimization (learning) problem (7) under the CP format (4), always has a minimizer $u^{\dagger}$, and this minimizer satisfies $\|u^{\dagger}-u^{*}\|_{H^{l}(\Omega)}\lesssim\varepsilon$ as $h\rightarrow 0$.

### v2

@@ Abstract
Machine learning solvers for partial differential equations (PDEs) have attracted growing interest. However, most existing approaches, such as neural network solvers, rely on stochastic training, which is inefficient and typically requires a great many training epochs. Gaussian process (GP)/kernel-based solvers, while mathematical principled, suffer from scalability issues when handling large numbers of collocation points often needed for challenging or higher-dimensional PDEs. To overcome these limitations, we propose TGPS, a tensor-GP-based solver that introduces factor functions along each input dimension using one-dimensional GPs and combines them via tensor decomposition to approximate the full solution. This design reduces the task to learning a collection of one-dimensional GPs, substantially lowering computational complexity, and enabling scalability to massive collocation sets. For efficient nonlinear PDE solving, we use a partial freezing strategy and Newton’s method to linerize the nonlinear terms. We then develop an alternating least squares (ALS) approach that admits closed-form updates, thereby substantially enhancing the training efficiency. We establish theoretical guarantees on the expressivity of our model, together with convergence proof and error analysis under standard regularity assumptions. Experiments on several benchmark PDEs demonstrate that our method achieves superior accuracy and efficiency compared to existing approaches. The code is released at https://github.com/BayesianAIGroup/TGPSolve-NonLinear-PDEs
@@ 1 Introduction
Machine learning (ML) solvers for partial differential equations (PDEs) have been receiving increasing attention due to their ease of implementation and competitive accuracy. These approaches approximate the solution with a machine learning model, such as deep neural networks (Raissi et al., 2019a), trained by minimizing a composite objective function that combines boundary and residual losses evaluated at a set of collocation points, thereby enforcing the boundary conditions and equation. Unlike traditional numerical solvers, ML approaches avoid complex, problem-specific discretization schemes and numerical routines, making them simpler and more convenient to implement and verify.
Despite these advantages, most existing ML solvers — including physics-informed neural networks (Raissi et al., 2019b) and recent Gaussian process (GP) and kernel-based methods (Fang et al., 2023; Xu et al., 2024) — rely on stochastic optimization to effectively learn the model parameters, which often requires tens of thousands to even millions of iterations, making solving procedure quite inefficient. GP and kernel-based solvers, though mathematically principled, also face scalability challenges as the number of collocation points grows. For example, Chen et al. (2021a); Long et al. (2022) place a GP prior over the solution and its derivatives, yielding block-structured covariance matrices whose time and memory costs exceed the standard $\mathcal{O}(M^{3})$ and $\mathcal{O}(M^{2})$ scaling, where $M$ is the number of collocation points. To mitigate this, Fang et al. (2023); Xu et al. (2025) proposed using product kernels on Cartesian grids, exploiting Kronecker algebra for efficient matrix operations. However, this approach requires estimating the solution values at all grid points, leading to exponential growth in parameters with dimension, and its reliance on structured grids limits applicability to irregular domains.
@@
Model: We introduce a set of one-dimensional factor functions, each modeled as a GP, along every input dimension. These factors are combined via multilinear tensor decompositions — such as CANDECOMP/PARAFAC (CP) (Harshman and others, 1970) or tensor-ring decomposition (Zhao et al., 2016) — to approximate the full solution. For each dimension, we place inducing points and represent the factor function by its GP conditional mean, with the inducing values serving as trainable parameters. Collocation points are freely sampled to form the training objective. This design allows our model to scale linearly with both the PDE dimension and the number of collocation points — not only in covariance matrix computation and storage but also in the number of trainable parameters.
@@ 2 Background
where ${\mathcal{F}}_{r}=\frac{1}{M_{\Omega}}\sum_{m=1}^{M_{\Omega}}\left({\mathcal{P}}(\text{NN}_{\Theta})({\bf x}_{m})-a({\bf x}_{m})\right)^{2}$ is the PDE residual loss, and ${\mathcal{F}}_{b}=\frac{1}{M-M_{\Omega}}\sum_{j=1}^{M-M_{\Omega}}\left({\mathcal{B}}\left(\text{NN}_{\Theta}\right)({\bf x}_{M_{\Omega}+j})-b({\bf x}_{M_{\Omega}+j})\right)^{2}$ is the boundary loss. Here NN denotes the neural network, $\Theta$ its parameters, and $\lambda_{b}>0$ a weighting coefficient. Training typically relies on stochastic optimization (e.g., ADAM (Kingma, 2014)), as in standard neural network applications like image classification. Achieving good accuracy generally requires tens of thousands of iterations, and a second-order optimizer (e.g., L-BFGS) is often used for refinement and stabilization (Shin, 2020; Li et al., 2023; Penwarden et al., 2023). As a result, while this framework is straightforward and convenient to implement, the training (solving) process is often lengthy and inefficient.
@@
Gaussian Process and Kernel-Based Solvers. An alternative class of solvers is based on GP/kernel methods, which rest on strong mathematical foundations. In (Chen et al., 2021a; Long et al., 2022) , a GP prior is placed over the solution $u$ and all its derivatives (or more generally, linear operators) $D_{j}(u)$ appearing in the PDE. The goal is to estimate the values of $u$ and all $D_{j}(u)$ evaluated at the collocation points, which leads to a multi-variate Gaussian prior distribution with a block covariance matrix, ${\bf C}=\{{\bf C}_{ij}\}$, where each block ${\bf C}_{ij}$ is associated with a pair of linear operators (e.g., derivatives and $u$ itself). Since each collocation point can contribute to multiple values (e.g., different $D_{j}(u)$), the covariance matrix $\bf C$} is typically larger than $M\times M$, where $M$ is the number of collocation points. As $M$ increases, the time and space complexity exceed $\mathcal{O}(M^{3})$ and $\mathcal{O}(M^{2})$, respectively, making computation prohibitively expensive or even infeasible for large $M$.
@@
To mitigate this issue, recent work by Xu et al. (2025) approximates the solution using standard GP/kernel interpolation: $u({\bf x};{\bf u}_{{\mathcal{M}}})=\kappa({\bf x},{\mathcal{M}}){\bf K}_{MM}^{-1}{\bf u}_{{\mathcal{M}}}$, where ${\bf K}_{MM}=\kappa({\mathcal{M}},{\mathcal{M}})$ is the $M\times M$ covariance matrix at the collocation points, and $\bf u$$_{{\mathcal{M}}}$ denotes the solution values at these points. Differential operators $D_{j}$ are applied directly to this interpolation to approximate $D_{j}(u)$, enabling the use of a training loss similar to (2). To reduce computation cost, collocation points are placed on a Cartesian grid: ${\mathcal{M}}={\bf s}^{1}\times\ldots\times{\bf s}^{d}$ where each ${\bf s}^{j}$ is a set of input locations along dimension $j$. By adopting a product kernel of the form $\kappa(\mathbf{x},\mathbf{x}^{\prime})=\prod_{j}\kappa_{j}(x_{j},x_{j}^{\prime})$, the covariance matrix admits a Kronecker product structure: $\kappa({\mathcal{M}},{\mathcal{M}})={\bf K}_{1}\otimes\ldots\otimes{\bf K}_{d}$, where each ${\bf K}_{j}=\kappa_{j}({\bf s}^{j},{\bf s}^{j})$ is the kernel matrix along input dimension $j$. Kronecker algebra (Kolda, 2006) allows one to avoid computing the full ${\bf K}_{MM}$; instead, only the smaller matrices ${\bf K}_{j}$ are computed and inverted, greatly reducing the cost. However, this method still requires estimating $\mathcal{U}$$_{{\mathcal{M}}}$ — the solution values over the entire grid $\mathcal{M}$}. As the input dimension increases, the size of $\mathcal{U}$$_{\mathcal{M}}$ grows exponentially, making storage and estimation infeasible. Furthermore, effective training still relies on stochastic optimization, often requiring up to one million iterations (Xu et al., 2025), rendering the solving procedure inefficient in practice.
@@ 3.1 Model
To leverage the principled mathematical framework of GPs/kernel methods while overcoming their scalability and efficiency bottlenecks, we propose TGPS, a tensor-GP based PDE solver. Specifically, we introduce a set of one-dimensional factor (component) functions for each input dimension $i$,
@@
where $1\leq i\leq d$, and ${\mathcal{G}}^{i}$ is a Reproducing Kernel Hilbert Space (RKHS) induced by a Mercer kernel function $\kappa_{i}(\cdot,\cdot)$. In other words, we model each factor function as a GP, $f^{i}_{r}\sim\mathcal{GP}(0,\kappa_{i}(\cdot,\cdot))$. For simplicity of discussion, we assume the PDE domain $\Omega\cup\partial\Omega\subseteq\bigtimes_{i=1}^{d}\Omega_{0}$. We then combine these factor functions via mutilinear tensor decomposition to construct the solution approximation. We consider two representative tensor decomposition models.
CANDECOMP/PARAFAC(CP) Decomposition (Harshman and others, 1970). We set $R_{1}=\ldots=R_{d}=R$, and model the solution function as
@@
Tensor-Ring (TR) Decomposition (Zhao et al., 2016). In each dimension $i$, we view the factor functions together as a single function with a matrix output, ${\bf F}^{i}=\{f^{i}_{r}(\cdot)\}:\mathbb{R}\rightarrow\mathbb{R}^{R_{i-1}\times R_{i}}$ where $R_{0}=R_{d}$. The solution is modeled as
@@ 3.2 Algorithm
Newton’s method (Tangent Linearization). Consider solving the PDE as a root finding problem: $\mathcal{R}(u)=0$, where $\mathcal{R}$} is the PDE operator. We apply Newton’s method by linearizing ${\mathcal{R}}(u)$ around the previous iteration $u^{\text{prev}}$ via a first-order Taylor expansion: ${\mathcal{R}}(u)\approx{\mathcal{R}}(u^{\text{prev}})+J(u^{\text{prev}})(u-u^{\text{prev}})=0$, where $J(u^{\text{prev}})$ is the Fréchet derivative, $J(u^{\text{prev}})=\frac{{\rm d}\mathcal{R}(u^{\text{prev}}+\epsilon v)}{{\rm d}\epsilon}\bigg|_{\epsilon=0}$, and $v$ is an arbitrarily small perturbation. By construction, this first-order approximation is always linear in $u$. In practice, the procedure amounts to replacing the nonlinear terms in the PDE and boundary conditions with their first-order Taylor expansions. For example, in (9), the cubic term $u^{3}$ is replaced by $(u^{\text{prev}})^{3}+3(u^{\text{prev}})^{2}(u-u^{\text{prev}})$.
@@ 4 Theoretical Analysis
Next, we establish the convergence result under the CP format. Our analysis follows the roadmap of Batlle et al. (2023); Xu et al. (2025), adopting standard assumptions on PDE stability and on the regularity of the domain and boundary conditions (Xu et al., 2025, Assumption 4.1).
@@
Let $u^{*}\in{\mathcal{U}}$ denote the unique strong solution of (1), and suppose Assumption 4.3 holds. Let ${\mathcal{M}}={\mathcal{M}}_{\Omega}\cup{\mathcal{M}}_{\partial\Omega}$ be a set of collocation points, with ${\mathcal{M}}_{\Omega}\subset\Omega$ and ${\mathcal{M}}_{\partial\Omega}\subset\partial\Omega$. Assume the Voronoi diagram induced by $\mathcal{M}$} has a uniformly bounded aspect ratio across all cells. Define the fill-distances $h_{\Omega}:=\underset{{\bf x}\in\Omega}{\sup}\;\underset{{\bf x}^{\prime}\in{\mathcal{M}}_{\Omega}}{\inf}|{\bf x}-{\bf x}^{\prime}|$ and $h_{\partial\Omega}:=\underset{{\bf x}\in\partial\Omega}{\sup}\;\;\underset{{\bf x}^{\prime}\in{\mathcal{M}}_{\partial\Omega}}{\inf}\rho_{\partial\Omega}({\bf x},{\bf x}^{\prime})$, where $|\cdot|$ is the Euclidean distance, and $\rho_{\partial\Omega}$ is a geodesic distance defined on $\partial\Omega$. Set $h=\text{max}(h_{\Omega},h_{\partial\Omega})$. Suppose each RKHS ${\mathcal{G}}^{i}$, to which the factor functions in dimension $i$ belong, is associated with a universal kernel, and that ${\mathcal{U}}={\mathcal{G}}^{1}\otimes\cdots\otimes{\mathcal{G}}^{d}$ [^1] Here $\mathcal{U}$} is a completed tensor product space, and $\otimes$ is the topological tensor product, i.e., the completion of the algebraic tensor product with respect to the product RKHS norm rather than the algebraic tensor product.. Then, for any arbitrarily small $\varepsilon>0$, with a sufficiently large $R$ and an appropriate $\delta$, the optimization (learning) problem (7) under the CP format (4), always has a minimizer $u^{\dagger}$, and this minimizer satisfies $\|u^{\dagger}-u^{*}\|_{H^{l}(\Omega)}\lesssim\varepsilon$ as $h\rightarrow 0$.